OpenStudy (anonymous):

given that cos(x)=1/3, find sin(90-x) please help!

geerky42 (geerky42):

\[\sin(90^o-x)=\cos(x)\]

OpenStudy (anonymous):

im not exactly sure where to start on the equation :(

geerky42 (geerky42):

you are already informed that \(\cos(x)=1/3\).

OpenStudy (aaronq):

Find x first, \(\sf cos^{-1}(cos(x))=cos^{-1}(\dfrac{1}{3})\) \(\sf x=cos^{-1}(\dfrac{1}{3})\)